# Charles Gaydon

### v3

@@
FRACTAL
Charles Gaydon
(April 2024)
[…]
FRACTAL: An Ultra-Large-Scale Aerial Lidar Dataset for 3D Semantic Segmentation of Diverse Landscapes
Charles Gaydon             Michel Daab             Floryne Roche
Institut national de l’information géographique et forestière (IGN), France
fractal.dataset@ign.fr
@@
AbstractMapping agencies are increasingly adopting Aerial Lidar Scanning (ALS) as a new tool to monitor territory and support public policies. Processing ALS data at scale requires efficient point classification methods that perform well over highly diverse territories. To evaluate them, researchers need large annotated Lidar datasets, however, current Lidar benchmark datasets have restricted scope and often cover a single urban area. To bridge this data gap, we present the FRench ALS Clouds from TArgeted Landscapes (FRACTAL) dataset: an ultra-large-scale aerial Lidar dataset made of 100,000 dense point clouds with high quality labels for 7 semantic classes and spanning 250 km². FRACTAL is built upon France’s nationwide open Lidar data. It achieves spatial and semantic diversity via a sampling scheme that explicitly concentrates rare classes and challenging landscapes from five French regions. It should support the development of 3D deep learning approaches for large-scale land monitoring. We describe the nature of the source data, the sampling workflow, the content of the resulting dataset, and provide an initial evaluation of segmentation performance using a performant 3D neural architecture.
Key Figures
➔ 250 km² of data sampled from 17,440 km² in 5 French regions
➔ 100,000 point clouds - 50 x 50 m each
➔ 9261M points from high density ALS (37 pts/m²)
➔ 7 semantic classes: other, ground, vegetation, building, water, bridge, permanent structure
I Introduction
I-A Aerial Lidar Scanning for land monitoring
@@
High-density ALS is now recognized as a powerful remote sensing modality to support important public actions such as ecological monitoring [1] and risk management (floods [2], forest fires [3]). Detailed 3D mapping also contribute to the consolidation of existing geographic databases. The reduction in costs associated with the acquisition, storage, processing, and dissemination of Aerial Lidar Scanning (ALS) data has paved the way for its production at unprecedented scales. In recent years, ALS has been increasingly used by public authorities at the regional and national levels. In a comprehensive report, [4] describe the availability of non-commercial Lidar data across Europe: at least a dozen European countries have already made nationwide Lidar acquisitions, with densities of 10 to 40 pts/m². In France, the national mapping agency (Institut national de l’information géographique et forestière, IGN) aims to map the entire French territory with high-density ALS point clouds (10 pulses/m², about 40 pts/m²) by 2026, in a program called Lidar HD (for "High Density") [5].
I-B Challenges and methods for point cloud classification
@@
Point cloud classification is a prerequisite to any downstream use of Lidar products. Ground points must be identified to produce accurate Digital Terrain Models (DTMs) and separated from vegetation points to normalize tree heights in forest applications, e.g. for forest biomass estimation. Buildings are reconstructed from building points for urban modeling, while city trees are mapped from clusters of vegetation points to manage urban biodiversity.
Understanding Lidar data requires specialized knowledge and tools. Many softwares exist for Lidar feature extraction (e.g. TopoDOT [6], TerraScan [7]…), however, each has its own strengths and weaknesses, and they all require some form of manual input to reach satisfying results [8]. The complexity, volume, and lack of structure of Lidar data make manual labelling time and labor-intensive. There is therefore a critical need for more automated methods.
Recently, there has been a surge in the development of deep learning approaches for the efficient semantic segmentation of point clouds. PointNet in 2016 [9] and PointNet++ in 2017 [10] paved the way by proposing a framework to process unordered point clouds without costly preprocessing. We refer the reader to Section 4 of a previous work [11] for a short history of the development of 3D deep learning architectures.
I-C The need for diversity in ALS benchmark datasets
@@
ALS differs from ground-based Lidar in critical ways: airborne Lidar has nadir orientation instead of lateral orientation, and a lower but more homogeneous point density [12]. Most importantly, airborne Lidar can be used to collect data at a much larger scale than ground-based Lidar, which adds specific challenges to their classification. Vast territories the size of a country have diverse landscapes and vegetation, along with unique signs of human activity, leading to high intraclass heterogeneity. Additionally, geographic data are characterized by spatial autocorrelation, and a long-tail distribution, resulting in a variety of rare scenes that may be spatially concentrated i.e. globally rare but locally frequent like wind turbines or greenhouses.
@@
The specific nature and scope of ALS must be addressed in the evaluation of 3D point cloud semantic segmentation methods. In doing so, having large representative annotated datasets is critical, however, current benchmark datasets come up short: they are limited in size and spatial diversity, typically covering a single area, and thus cannot attest to the capacity of 3D deep learning models to deal with the challenges of large-scale land monitoring.
[13] make a comprehensive review of current Lidar benchmark datasets for 3D point cloud semantic segmentation as of October 2023. They list 26 datasets of which only 6 have ALS data, which we report in Table I. To date, OpenGF [14] is the largest of them: it spans almost 50 km² and explicitly includes diverse landscapes. Interestingly, OpenGF is built from open large-scale ALS archives from 4 countries in 3 continents, via a careful selection of areas with high-quality point annotations in diverse landscapes (i.e. metropolis, small city, village, mountain). Authors demonstrate the effectiveness of leveraging open Lidar assets to quickly create high-quality, diverse, large-scale benchmark datasets for 3D semantic classification.
OpenGF is, however, tailored for the specific task of ground/non-ground classification and was stripped of the semantic classes used in most applications (e.g., building, vegetation…). The other ALS benchmark datasets only span a single urban area, and we observe that rich semantics and high point densities are only made available in small volumes. The Dayton Annotated Lidar Earth Scan (DALES) [12] attempts to offer the level of detail required for typical urban applications, with 8 semantic classes and a high point density of 50 pts/m². Covering an area of 10 km², DALES is the largest dataset of its kind. However, with a single urban area, it falls far short of the landscape diversity found in EO applications.
@@
TABLE I: Benchmark datasets for 3D semantic segmentation of ALS data. We include the datasets listed by [13] along with their proposed benchmark dataset CENAGIS-ALS, and our own dataset FRACTAL. ALS: Aerial Lidar Scanning; ULS: Unmanned Lidar Scanning.
@@
I-D Contributions
This data gap motivates us to release the FRench ALS Clouds from TArgeted Landscapes, or FRACTAL. FRACTAL is a large dataset for the semantic segmentation of ALS data that spans 250 km². FRACTAL is made of 100,000 distinct point clouds, each spanning 50 $\times$ 50 m and typically containing 20k to 40k labelled points (37 pts/m² on average).
The dataset is built from the French Lidar HD data [5], using a sampling scheme that explicitly concentrates rare classes, rare objects, and challenging landscapes. Being sampled from an initial area of 17,440 km² in 5 French regions, it is characterized by a high diversity of scenes.
Point clouds are labelled with seven common semantic classes: ground, vegetation, building, water, bridge, permanent structure, and other. The labels were produced with automated processes and then verified and corrected by Lidar operators to achieve high quality classification. FRACTAL is overall the largest open benchmark dataset for 3D semantic segmentation, and the first to offer the diversity of landscapes inherent to large-scale land monitoring. It provides a challenging benchmark dedicated to the development of 3D deep learning methods.
We make the following contributions:
  • Set a baseline evaluation of segmentation performance on the dataset using a state-of-the-art deep learning method to show its potential for benchmarking.
Section II presents the data sources used to create FRACTAL. Section III describes the strategy used for cataloguing and sampling data patches at scale into a deep learning benchmark dataset. Section IV presents the resulting dataset itself.
II Data Sources
@@ II-A ALS point clouds
The Lidar HD program [5] is a national initiative that aims to provide a 3D description of the French territory by 2026, using high-density ALS (10 pulses/m² or about 40 pts/m²). Designed to meet the needs of public actions (flood prevention, estimation of forest resources, etc.), it is operated by the IGN with both internal resources and subcontracted assistance. The data acquired and produced as part of this program (i.e., point clouds, Digital Terrain Models, Digital Surface Models) are made available as open data with extensive documentation [19].
@@
The program covers mainland France and its overseas territories for a total of 550,000 km². The program consists of four phases: data acquisition, storing, processing, and dissemination, along user support. The data acquisition and processing are sequenced in blocks of 50 x 50 km and must be compatible with a variety of Lidar sensors (Leica, Riegl, and Teledyne/Optech), acquisition seasons (leaf-on or leaf-off), and landscapes (urban, rural, mountains, seashores, overseas territories, etc.).
Processing includes the semantic segmentation of point clouds into 11 classes: unclassified, ground, vegetation (low, medium, and high), buildings, water, bridge deck, permanent structures, artefact, synthetic. We detail the specification of this nomenclature in Appendix LABEL:appendix:lhd_nomenclature.
In addition, the classification of Lidar HD data may come in one of two flavors:
@@
Classified Lidar HD: results from a fully automated classification process operated by the IGN, consisting of explicit detection methods using the IGN’s national databases and the results of AI deep learning models.
@@
Optimized Lidar HD: results from automatic classification processing followed with manual corrections by subcontractors. The classification is audited by the IGN via a visual inspection of 10% of the delivered point clouds. In the audit, all classification errors are listed and rated for severity, with particular attention to confusions in buildings, bridges and ground. Consolidated error scores are computed, first for each of these three critical semantic classes, then globally. These scores are required to be under specific thresholds for the classification to be validated. Particular attention is given to the exhaustivity of individual buildings, with a target recall of 99.9% in high-stakes areas (e.g. flood-prone, urban…) and a target recall of 99.5% elsewhere.
II-B VHR aerial images
@@
The ORTHO HR® [20] is a mosaic of VHR aerial images acquired during national surveys. Every year, about a third of French metropolitan territory is updated. The individual images are mapped onto a cartographic coordinate reference system and projected on the RGE ALTI Digital Terrain Model for orthorectification. ORTHO HR images have a high spatial resolution of 0.20 m, and near infrared, red, green and blue channels. Radiometric processing methods, including equalization and global correction, are applied to obtain the final product. The radiometry of the red, green, and blue channels cannot be considered a physical measurement of channel reflectance due to variations in acquisition dates and specific radiometric corrections applied for visual appealing.
@@
Aerial surveys are not synchronized with Lidar HD acquisitions and a variable time lag of up to 3 years might separate them. This is because aerial images from the ORTHO HR are renewed every three years, while Lidar data are processed as soon as they are acquired. In the interim, buildings may have been constructed and vegetation may have expanded or been cleared. Moving elements such as vehicles are unlikely to be consistent. Also, the appearance of cultivated fields is likely to be inconsistent between modalities due to different seasons of acquisition.
III Sampling Methodology
III-A Sampling area
@@
At the time of dataset creation, Lidar HD data is available for about half of metropolitan France, mainly in its southern half. To have the highest possible annotation quality, we restrict our sampling to Optimized Lidar HD i.e., to the areas that went through human verification and audit of the classification. Figure 1 presents the 5 spatial domains we consider, each spanning 3456 km² on average, and distant from at least 100 km from one another. The spatial domains compose a sampling area of 17,440 km² in Southern France.
@@
To define a common setting for the benchmark of deep learning models, we reserve an area for testing in each spatial domain. Each of the 5 test areas span 210 km² on average, for a total test sampling area of 1049 km². Having only a limited number of spatial domains with their own unique characteristics (seashores, mountains, field crops, etc.), test data and train data share the same spatial domains. Test areas are however contiguous, large, and distinct from the train areas. This enables spatial block validation and reduces the influence of spatial autocorrelation on model evaluation, which is critical when evaluating remote sensing solutions [21].
@@
Figure 1: The five spatial domains composing the 17,440 km² of the sampling area. The 1,049 km² of area reserved to sample the test set are highlighted in red.
III-B Dataset size and sample unit
We aim to produce a manageable dataset with suitable size for common research practices and computing infrastructures. Based on our own experience with training semantic segmentation models on Lidar HD data, we set a target dataset size of 250 km² - a 70-fold reduction compared to the initial sampling area.
@@ III-B Dataset size and sample unit
The sample unit is a 50 × 50 m square patch. We have found that such a width provides a good balance between detailed structures and long-range dependencies. A square shape facilitates data extraction and visualization of data patches. Also, training models on square point clouds favors the use of non-overlapping tiling at inference time, thus gaining efficiency by limiting redundant predictions.
III-C Considerations for dataset diversity
We adopt the following design principles to sample the dataset:
@@ III-C Considerations for dataset diversity
Class rebalancing: We explicitly address class imbalance, which is often detrimental to model training, by up-sampling rarer classes such as water, bridge, and permanent structure.
@@
Uncertainty sampling: We aim to increase the representation of scenes known to be challenging for classification models by upsampling specific landscapes and hardscapes, such as mountainous areas, seashores, or complex urban scenes.
@@
Spatial sampling as a proxy for scene diversity: Given the spatial autocorrelation of geographic data, we aim for the broadest spatial distribution of samples. This maximizes diversity within each category of scenes and globally.
III-D Cataloguing of available data
For efficient consideration of classes and landscapes, we need a detailed representation of all available data of the sampling area with patch resolution. This representation takes the form of a geographical database called the Lidar Patch Catalog (LiPaC).
Cataloging data boils down to two steps, listing and describing, which we illustrate in Figure 2. We start by regularly dividing the sampling area into $50\times 50$ m patches. We then describe all patches with two sets of descriptors: Lidar descriptors and vector descriptors. Lidar descriptors are derived from the Lidar data patches themselves: number of points in each semantic class, elevation (above sea level), and elevation gain. Vector descriptors are derived by spatial crossing with the BD TOPO® [22], a 3D vector description of the French territory and its infrastructures (buildings, roads, etc.). In the LiPaC, we include information about the following classes of objects: greenhouses, antennas, pylons, highways, parking lots, water surfaces, greenhouses, pylons, antennas, and industrial, commercial, or agricultural buildings.
Figure 2: Key steps involved in setting up a Lidar Patch Catalog (LiPaC) i.e. the geographical database that we use to a) list and b) describe all candidate Lidar patches for sampling.
The LiPaC is a comprehensive tool for flexible and efficient sampling of all Lidar tiles under consideration. Its implementation as a PostGIS database gives us the full power of SQL syntax to select data and further characterize each Lidar patch with bespoke SQL queries that combine simple descriptors into more complex ones. We define two categories of high-level Boolean flags, which we report in Figure II. The first category contains 12 scene flags related to the semantic classification of the point cloud and the presence of certain objects within a semantic class. For example, we define a BUILD flag for patches that contain some points classified as buildings (n $\geq$ 500 pts) and a BUILD_GREENHOUSE flag for patches that intersect greenhouses. The second category contains 7 scene flags that characterize patches in terms of landscapes, using heuristic definitions. For example, a patch is flagged as SEASHORE if it contains water (n $\geq$ 50 pts) and ground (n $\geq$ 100 pts) and is within 10 m of sea level.
@@ III-D Cataloguing of available data
TABLE II: Scene types and their target minimal concentration for targeted sampling.
@@
III-E Patch sampling
Having a detailed description of each patch makes it easy to rebalance semantic classes and ensure that challenging or rarer types of scenes are represented in the dataset, following the design principles established in Section III-C. We adopt a simple but effective sampling scheme consisting of targeted sampling followed by completion sampling:
Targeted sampling: For each scene type (i.e., flag), we perform spatially stratified sampling, i.e., we sample patches with stratification on 1 $\times$ 1 km Lidar HD tiles until a minimum concentration of that scene type is reached. Each sampling is performed independently, with replacement, to ensure that the spatial diversity within each group is optimal. As a result, the final proportion of each scene type may be higher than the minimum requirement. We set a target requirement for each scene type so that approximately 50% of the final dataset is found using targeted sampling (see Table II).
Completion sampling: The rest of the dataset is found by spatially stratified sampling without considering the flags. This is to ensure that we include more ordinary scenes such as field crops and rural areas, which are the majority.
III-F Dataset split
We define a reference split into train, val, and test sets with an 80/10/10 ratio: 225 km² of data is used for model training, of which 25 km² is reserved for in-training evaluation, and 25 km² of data is sampled from test areas and kept for model evaluation. Train and test areas are sampled independently with the same sampling parameters.
Patches for in-training validation are sampled from the train areas with spatially stratified stratification, independently for each scene type. This results in the widest possible spatial distribution for each type of scene and ensures that all landscapes are equally represented in the train and val sets. Refer to Appendix LABEL:appendix:descriptors_trainvaltest for the proportions of scene types in each set.
III-G Data extraction and colorization
Point clouds are made available in the LAZ 1.4 format defined by the American Society for Photogrammetry and Remote Sensing (ASPRS) [23]. The following dimensions are particularly relevant for semantic segmentation: xyz coordinates in Lambert 93 spatial reference system (EPSG:2154), intensity (16-bits encoding), return number and number of returns, scan angle, and near-infrared, red, green and blue (16-bits encoding). The Lidar HD classification nomenclature, described in Appendix LABEL:appendix:lhd_nomenclature, is left intact in the data patches.
IV FRACTAL: The Dataset
In this subsection, we present the resulting dataset, which we name FRACTAL: FRench ALS Clouds from TArgeted Landscapes.
Section IV-A presents the characteristics of the dataset itself and emphasizes its diversity. Section IV-B documents the file structure itself. Section IV-C list limitations to consider when using the dataset.
IV-A Spatial and semantic diversity
Table III presents key figures of the dataset, namely its size expressed in number of points, number of tiles, number of patches, and total area. We compare these figures with those of the sampling area to highlight that, while the dataset has been substantially reduced in size to facilitate manageable research, it still retains spatial diversity. Despite being 70 times smaller, the dataset incorporates data from all 17,440 initial tiles, ensuring comprehensive representation.
TABLE III: Size, areas, and number of points in FRACTAL compared to the sampling area. $r$ = ratio.
Sampling area $\rightarrow$ FRACTAL $\mathbf{r}$

### v4

# FRACTAL: An Ultra-Large-Scale Aerial Lidar Dataset for 3D Semantic Segmentation of Diverse Landscapes

@@
Abstract
@@
Mapping agencies are increasingly adopting Aerial Lidar Scanning (ALS) as a new tool to map buildings and other above-ground structures. Processing ALS data at scale requires efficient point classification methods that perform well over highly diverse territories. Large annotated Lidar datasets are needed to evaluate these classification methods, however, current Lidar benchmarks have restricted scope and often cover a single urban area. To bridge this data gap, we introduce the FRench ALS Clouds from TArgeted Landscapes (FRACTAL) dataset: an ultra-large-scale aerial Lidar dataset made of 100,000 dense point clouds with high quality labels for 7 semantic classes and spanning 250 km². FRACTAL achieves high spatial and semantic diversity by explicitly sampling rare classes and challenging landscapes from five different regions of France. We describe the data collection, annotation, and curation process of the dataset. We provide baseline semantic segmentation results using a state of the art 3D point cloud classification model. FRACTAL aims to support the development of 3D deep learning approaches for large-scale land monitoring.
1 Introduction
@@
High-density Aerial Lidar Scanning (ALS) is now recognized as a powerful remote sensing modality to support important public actions such as ecological monitoring [17] and risk management (e.g., for floods [18] and forest fires [26] prevention). Detailed 3D mapping also contribute to the consolidation of existing geographic databases (e.g., by updating building footprints). The reduction in costs associated with the acquisition, storage, processing, and dissemination of Aerial Lidar Scanning (ALS) data has paved the way for its production at unprecedented scales.
Figure 1: Random scenes from our FRACTAL dataset. Each point is colored according to its semantic class; ground (orange), vegetation (greens), building (red), water (cyan), bridge (yellow), permanent structure (purple), other (white).
In recent years, ALS has been increasingly used by public authorities at the regional and national levels. In a comprehensive report, [14] describe the availability of non-commercial Lidar data across Europe: at least a dozen European countries have already performed nationwide Lidar acquisitions, with densities from 10 to 40 pts/m². In France, the national mapping agency (Institut national de l’information géographique et forestière, IGN) aims to map the entire French territory with high-density ALS point clouds (10 pulses/m², about 40 pts/m²) by 2026, in a program called Lidar HD (for ”High Density”) [10].
@@
Point cloud classification is a prerequisite to most downstream use of Lidar products. For example, ground points must be identified to produce accurate Digital Terrain Models (DTMs) and separated from vegetation points to normalize tree heights in forest applications, e.g., for forest biomass estimation. Buildings are reconstructed from building points for urban modeling, while city trees are mapped from clusters of vegetation points to manage urban biodiversity.
The complexity, volume, and lack of structure of Lidar data make manual labelling of Lidar data time and labor-intensive; existing commercial software all require some form of manual input to reach satisfying results [27]. Deep learning methods are increasingly available, especially since the publication of PointNet [20] in 2016, and of PointNet++ [21] a year later.
Researchers need large annotated Lidar datasets to develop and evaluate their 3D deep learning methods. This need motivates us to release the FRench ALS Clouds from TArgeted Landscapes, or FRACTAL. FRACTAL is a large dataset for the semantic segmentation of ALS data, that spans 250 km². FRACTAL is made of 100,000 distinct point clouds, each spanning 50 $\times$ 50 m and typically containing 20k to 40k labelled points (37 pts/m² on average).
The dataset is built from the French Lidar HD data [10], using a sampling scheme that explicitly concentrates rare classes, rare objects, and challenging landscapes. Being sampled from an initial area of 17,440 km² in 5 French regions, it is characterized by a high variety of scenes.
Point clouds are labelled with seven common semantic classes: ground, vegetation, building, water, bridge, permanent structure, and other. The labels were produced with automated processes and then verified and corrected by Lidar operators to achieve high quality classification. FRACTAL is overall the largest open benchmark dataset for 3D semantic segmentation, and the first to offer the diversity of landscapes inherent to large-scale land monitoring.
We make the following contributions:
  • Set a baseline evaluation of segmentation performance on the dataset using a state-of-the-art deep point cloud segmentation model, to show its potential for benchmarking.
Section 2 lays out current ALS benchmarks and their shortcomings. Section 3 introduces the data sources used to create FRACTAL. Section 4 presents the strategy for cataloguing and sampling data patches at scale, and highlights the high diversity of the resulting dataset. Section 5 describes the first experimental results on the benchmark.
2 Related work
@@
ALS differs from ground-based Lidar in critical ways: airborne Lidar has nadir orientation instead of lateral orientation, and a lower but more homogeneous point density [24]. Most importantly, airborne Lidar can be used to collect data at a much larger scale than ground-based Lidar, which adds specific challenges to their classification. Vast territories the size of a country have diverse landscapes and vegetation, along with unique signs of human activity, leading to high intraclass heterogeneity. Additionally, geographic data are characterized by spatial autocorrelation and a long-tail distribution, resulting in a variety of rare scenes that may be spatially concentrated, i.e., globally rare but locally frequent, like wind turbines or greenhouses.
@@
The specific nature and scope of ALS must be addressed in the evaluation of 3D point cloud semantic segmentation methods. Having large representative annotated datasets is critical, however, current benchmark datasets come up short: limited in size and spatial diversity, they cannot attest to the capacity of 3D deep learning models to deal with the challenges of large-scale land monitoring. In a comprehensive review of current Lidar benchmark datasets for 3D point cloud semantic segmentation, authors found 26 datasets [28]; only six of them have ALS data, which we report in Tab. 1 along with their proposed dataset (CENAGIS-ALS) and ours (FRACTAL).
@@
Table 1: Benchmark datasets for 3D semantic segmentation of ALS data. We include the datasets listed by [28] along with their proposed benchmark dataset CENAGIS-ALS, and our own dataset FRACTAL. ALS: Aerial Lidar Scanning; ULS: Unmanned Lidar Scanning.
@@
Five of these benchmarks cover only a tiny area, not exceeding 2 km². This reduced scale contrasts with the larger support that other modalities enjoy in land monitoring, e.g., with massive land cover mapping benchmarks such as LandCover.ai [2], which covers 216 km² with VHR aerial images, and FLAIR [6], which spans 817 km² across 50 distinct spatial domains with both VHR aerial images and Sentinel 2 time series.
Furthermore, point densities in these ALS benchmarks is either too low for typical urban applications (3-4 pts/m² in LASDU [25], 4-7 pts/m² in ISPRS Vaihingen [19]) or unreasonably high (275 pts/m² in CENAGIS-ALS [28], 348 pts/m² in DublinCity [29], 800 pts/m² in Hessigheim 3D [16]).
The Dayton Annotated Lidar Earth Scan (DALES) [24] attempts to provide a level of detail more representative of real-world use cases, with 8 semantic classes and a point density of 50 pts/m². It is a significant improvement in scale compared to previous urban ALS benchmarks: covering an area of 10 km², DALES is the largest dataset of its kind.
Unfortunately, all ALS benchmarks mentioned so far, including DALES, are limited to single urban areas. This results in adjacent training and test data, which is likely to result in overestimated performance metrics given the high autocorrelation of geographic data. Even with reduced ambition (e.g., only considering urban contexts), it is impossible to draw conclusions about the spatial generalizability of the benchmarked methods with current ALS benchmarks. In addition, considering only a single, small urban area hides the complexity of regional scale land monitoring. Overall, researchers need datasets with better territorial representativeness.
OpenGF [22] is the one ALS benchmark that addresses these shortcomings. It is built from open large-scale ALS archives from 4 countries in 3 continents, via a careful selection of areas with high-quality point annotations in diverse landscapes (i.e., metropolis, small city, village, mountain). OpenGF is largest ALS benchmark to date, with 542M points over 50 km². Authors demonstrate the benefits of leveraging open Lidar assets to quickly create high-quality, diverse, large-scale benchmark datasets for 3D semantic classification. However, OpenGF is tailored for the specific task of ground/non-ground classification and was stripped of the semantic classes used in most applications (e.g., building, vegetation).
In summary, current ALS benchmarks for 3D point cloud segmentation are either limited in both volume and representativeness, and unsuitable to evaluate generalizability of methods, or are large and diverse enough but lack semantic depth (OpenGF). Our proposed dataset, FRACTAL, aims to address these shortcomings with: volume (250 km²), territorial representativeness (sampling from 17,440 km² in 5 distinct spatial domains), reduced influence of spatial autocorrelation (large contiguous test areas), and a semantic depth matching typical land monitoring applications (7 classes).
3 Data Collection
@@ 3.1 ALS point clouds
The Lidar HD program [10] is a national initiative that aims to provide a 3D description of the French territory by 2026, using high-density ALS (10 pulses/m² or about 40 pts/m²). The data produced as part of this program (i.e., point clouds, Digital Terrain Models, Digital Surface Models) are made available as open data with extensive documentation [11].
@@
The program covers mainland France and its overseas territories for a total of 550,000 km². It consists of four phases: data acquisition, storing, processing, and dissemination, along user support. The data acquisition and processing are sequenced in blocks of 50 $\times$ 50 km and must be compatible with a variety of Lidar sensors (Leica, Riegl, and Teledyne/Optech), acquisition seasons (leaf-on or leaf-off), and landscapes (e.g., urban, rural, mountains, seashores, overseas territories). The point clouds are disseminated with semantic segmentation labels from 11 classes: unclassified, ground, vegetation (low, medium, and high), buildings, water, bridge deck, permanent structures, artifact, synthetic. The specification of this nomenclature is detailed in Sec. D.1.
The classification of Lidar HD data comes in two flavors:
@@
Classified Lidar HD   Results from a fully automated classification process using commercial software and deep learning models.
@@
Optimized Lidar HD   Results from automatic classification followed by manual corrections. The annotations are then audited via visual inspection for approximately 10% of the delivered point clouds. All classification errors are listed and rated for severity, with particular attention to confusions in buildings, bridges and ground. Special importance is given to the exhaustivity of individual buildings, with a minimal recall of 99.9% in high-stakes areas (e.g., flood-prone, urban) and a minimal recall of 99.5% elsewhere. The classification is validated if it meets the requirements for real-world use; otherwise, it undergoes further manual corrections until it reaches the desired quality.
3.2 VHR aerial images
@@
The ORTHO HR® [12] is a mosaic of VHR aerial images acquired during national surveys. The individual images are mapped onto a cartographic coordinate reference system and projected on the RGE ALTI Digital Terrain Model for orthorectification. ORTHO HR images have a high spatial resolution of 0.20 m, and near infrared, red, green, and blue channels. Radiometric processing methods, including equalization and global correction, are applied to obtain the final product.
@@
Aerial surveys are not synchronized with Lidar HD acquisitions and a variable time lag of up to 3 years might separate them. Indeed, the aerial images are renewed every three years, whilst Lidar data are processed as soon as they are acquired. In the interim, buildings may have been constructed and vegetation may have expanded or been cleared. Moving elements such as vehicles are unlikely to be consistent. Also, the appearance of cultivated fields is likely to be inconsistent between modalities due to different seasons of acquisition. Despite these discrepancies, working with colorized point clouds is recommended as it systematically yields better segmentation models. For all practical purposes, we document the year of acquisition of images and point clouds as part of the dataset’s metadata.
4 FRACTAL: the dataset
4.1 Area of interest
@@
At the time of dataset creation, Lidar HD data is available for about half of metropolitan France, mainly in its southern half. To have the highest possible annotation quality, we restrict our sampling to Optimized Lidar HD i.e., to the areas that went through human verification and audit of the classification. Fig. 2 presents the 5 spatial domains we consider, each spanning 3456 km² on average, and at least 100 km distant from one another. The spatial domains compose an area of interest of 17,440 km² in Southern France, acquired with a variety of Lidar sensor (documented in Appendix D.3).
@@
To define a common setting for the benchmark of deep learning models, we reserve an area for testing in each spatial domain. Each of the 5 test areas span 210 km² on average, for a total of 1049 km². Having only a limited number of spatial domains with their own unique characteristics (seashores, mountains, field crops, etc.), test data and train data share the same spatial domains. Test areas are however contiguous, large, and distinct from the train areas. This enables spatial block validation and reduces the influence of spatial autocorrelation on model evaluation, which is critical when evaluating remote sensing solutions [15].
@@
Figure 2: The five spatial domains composing the 17,440 km² of the area of interest. The 1,049 km² reserved to sample the test set are highlighted in red.
4.2 Sampling methodology
@@
The sample unit is a 50 × 50 m square patch, which provides a compromise between detailed structures and long-range spatial dependencies. A square shape facilitates data extraction and visualization of data patches. Also, training models on square point clouds favors the use of non-overlapping tiling at inference time, thus gaining efficiency by limiting redundant predictions.
We adopt the following design principles for sampling:
@@
Class rebalancing: We explicitly address class imbalance, which is often detrimental to model training, by oversampling rarer classes such as water, bridge, and permanent structures.
@@
Uncertainty sampling: We aim to increase the representation of scenes known to be challenging for classification models by oversampling specific landscapes and hardscapes, such as mountainous areas, seashores, or complex urban scenes.
@@
Spatial sampling as a proxy for landscape diversity: Given the spatial autocorrelation of geographic data, we aim for the broadest spatial distribution of patches. This maximizes diversity within each category of scenes, and globally.
Scenes of special interest need first to be identified as such to consider them for sampling. We define 19 scene types, which we list in Tab. 2. Twelve of them are related to the semantic classification of the point cloud and to the presence of certain objects within a semantic class. The seven other characterize patches in terms of landscapes, using heuristic definitions.
@@
Table 2: Scene types and their target minimal proportion for targeted sampling.
@@
We adopt a simple yet effective sampling scheme consisting of targeted sampling followed by completion sampling:
Targeted sampling: For each scene type, we sample patches with stratification on 1 $\times$ 1 km Lidar HD tiles, until we reach the target proportion for that scene type. Each sampling is performed independently, with replacement, to ensure that the spatial diversity within each group is optimal; duplicates are then dropped.
Completion sampling: The dataset is completed with stratification on 1 $\times$ 1 km Lidar HD tiles, this time without consideration for scene type. This is to ensure that we include more ordinary scenes such as field crops and rural areas, which are the majority.
Based on our own experience with training semantic segmentation models on Lidar HD data, we set a target dataset size of 250 km², or 100,000 patches – a 70-fold reduction compared to the initial area of interest. This manageable size is suitable for common research practices and computing infrastructures.
For model benchmarking, we define a reference train/val/test split with an 80/10/10 ratio: 225 km² of data go to model training, of which 25 km² are reserved for in-training evaluation, and an additional 25 km² of data are sampled from test areas and kept for model evaluation. Train and test areas are sampled independently with the same target proportions of each scene type. Patches for in-training validation are sampled from the train areas, again with stratification on Lidar HD tiles. This strategy results in the widest possible spatial distribution for each scene type, and ensures that all landscapes are equally represented in the train, val, and test sets. See Appendix B.2 for the proportions of scene types in each set.
4.3 Spatial and landscape diversity in FRACTAL
Tab. 3 presents key figures of the dataset. Despite being 70 times smaller than the area of interest, the dataset incorporates data from all 17,440 initial Lidar HD tiles, ensuring comprehensive representation of its landscapes. With 9,261M points in 100,000 patches scattered over five vast spatial domains, FRACTAL has unprecedented territorial representativity, spatial diversity, and volume.
Table 3: Size, areas, and number of points in FRACTAL compared to the area of interest. $r$ = ratio.
Area of interest $\rightarrow$ FRACTAL $\mathbf{r}$